OpenStudy (apoorvk):

Determine all such integers a and b for which one of the roots of 3x^3+ax^2+bx+12=0 is equal to 1 + 3^(1/2).

OpenStudy (anonymous):

if 1+sqrt3 is a root then so is 1-sqrt3: (x-(1+sqrt3))(x-(1-sqrt3))*P(x) = 0 (x^2 - 2x -2)*P(x) = 0 P(x) must be linear in the form: (mx + n). So now you can mutiply out: (x^2 -2x - 2)*(mx + n). now equate the two expressions coeficient to coeficient an I got a = -12, b = 6.

OpenStudy (apoorvk):

thanks buddy.... can you also check out my previous question??

OpenStudy (anonymous):

where?

OpenStudy (apoorvk):

on my profile page